) AN ELECTRON JUMPS FROM 4th ENERGY LEVEL TO 1st ENERGY LEVEL IN AN HYDROGEN ATOM. HOW MANY PHOTONS OF ENERGY ARE EMITTED ?

Grade:11

1 Answers

rutvik
8 years ago
v = Rz[ 1/x2 – 1/y]
   z=1 for h-atom ( becoz.. x
  x=1 and y=4
  1/\lambda = 109678/cm [ 1/12-1/42 ]
         = 109678 * 15/16
     \lambda= 16/109678 * 15cm
        = 0.000009725
        = 9725 * 10-9 cm
        = 9725 * 10-11 m